OpenStudy (raimondas):

how to continue? lg (3^x-3) -lg2=lg (3^x+9)-lg(3^x-3) 3^x-3/2=3^x+9/3^x-3 (3^x-3)^2 =2(3^x+9) 3^x=a a^2-6a+9=2a+18 a^2-8a-9=0 D=100 (sqr)D=10 x1=-2 x2=8 3^x=-2 3^x=8 how to continue solving? or is there any mistakes ? need to find x that will be correct to inequality

hartnn (hartnn):

i do see one error. a^2 -8a -9 =0 a^2-9a +a -9 = 0 a(a-9) + 1 (a-9) = 0 (a+1)(a-9) = 0 a=-1,9 3^x = -1,9 lets continue, 3^x cannot be negative as the least value is 0. so discarding 3^x = -1 3^x = 9 we know 3^2 = 9 x=2 let me know if you have any more doubts :)

hartnn (hartnn):

no doubts? :)`